OpenStudy (anonymous):

ln[\z+2\]-ln[\z+3\]=ln[\t\]+c then, i apply e^ on both side e^(ln[\(z+2)/(z+3)\]=e^(ln[\t\]+c) =[\(z+2)/(z+3)\] = [\(t)(e^c)\] where e^c =k [\(z+2)/(z+3)\]=[\(k t)\] is this right??? or e^(ln[(z+2)/(z+3)] =e^(ln t)+e^c then (z+2)/(z+3)=t+K where K=e^c which one is right??

OpenStudy (experimentx):

\[ at the beginning

OpenStudy (anonymous):

you mean first one is right?

OpenStudy (experimentx):

for latex ... use \[ at the beginning

OpenStudy (anonymous):

oops. thank you !

zepdrix (zepdrix):

Yes, the first method you applied was correct. :)

OpenStudy (anonymous):

thank you! can u explain why is like that?? i mean why i cant e to the each lnt and c

zepdrix (zepdrix):

You have to exponentiate the `entire side`. Which gives,\[\Large e^{\left(\ln t+c\right)}\]A handy rule of exponents tells us:\[\Large a^{b+c} \quad=\quad a^b\cdot a^c\]

zepdrix (zepdrix):

\[\Large e^{\left(\ln t+c\right)} \quad=\quad e^{\ln t}\cdot e^{c}\]

zepdrix (zepdrix):

Why can't you exponentiate them individually? I'm not sure why.. I don't know if there exists an operation that allows us to do that :o hmm

OpenStudy (anonymous):

thank you!!! i was just arguing about this with gf , i was keep telling her have to exp whole part. but just cant explain why..

zepdrix (zepdrix):

The same applies in reverse. If you had something like,\[\Large stuff\quad=\quad e^t+C\]You wouldn't be able to apply the log to each term individually.\[\Large \ln(stuff) \quad=\quad \ln(e^t+C)\]

zepdrix (zepdrix):

In ^ that particular example you would probably want to move the C to the other side so you could do sneaky stuff to the e^t :)\[\Large stuff-C=e^t\]But whatev, it's just an example ^^ heh

OpenStudy (anonymous):

thank you very much, very clear explanation.